Question Part
Points
Model the data using an exponential function f(x) = Abx. HINT [See Example 1.]
x 0 1 2
f(x) 600 300 150
f(x) =

Answer by Boreal(15235)

Exponential function is ab^x
600=a*b^0=a
300=ab^1
b=300/a
b=0.5
function should be 600*(0.5)^x ANSWER
x=0, function is 600
x=1, function is 300
x=2 function is 600*0.5^2=150